OpenStudy (anonymous):

f(x) = I x-2 I x= -1

hero (hero):

f(x) = |x - 2| f(-1) = |(-1) -2|

OpenStudy (anonymous):

-3?

hero (hero):

(-1) - 2 = -3 |(-1) - 2| = |-3| = ?

OpenStudy (anonymous):

so it would have to be a positive number... right... so the answer is 3?

hero (hero):

Good job

OpenStudy (anonymous):

Ok, so if th "X" value is 0, then the answer would be... 2

hero (hero):

Yes, correct